13. Ammonia, NHs, is produced industrially from nitrogen and hydrogen as follows: N2(g)+ 3H2(g) 0 2NHs(g) What mass, of which starting material, will remain when 30.0 g of N2 and 10.0 g of H2 react untilthe limiting reagent is completely consumed?

help please
Answer #1

N2(g)  +  3 H2(g)  eqarrow.png  2 NH3(g)

1mole N2 react with 3 moles of H2 gives 2 moles of NH3

mass relation: 28 g N2 react with 6 g of H2

30 g N2 needs 30x6/28 = 6.43g H2

so H2 is excess reagent

Excess reagent remains = 10-6,43 = 3.57g